OpenStudy (anonymous):

Solve -1/8c=2 A.-4 B.16 C.-16 D.4

OpenStudy (ageta):

Let's solve your equation step-by-step. −1 8 c=2 Step 1: Divide both sides by (-1)/8. −1 8 c −1 8 = 2 −1 8 c=−16 Answer: c=−16

OpenStudy (anonymous):

c=-16

OpenStudy (anonymous):

-1/8(-16)=2

OpenStudy (ageta):

tnxs